# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ 1 Introduction
To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors to model temporal tissue dynamics (Figure 1). By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories. Table 1 highlights the key advantages of ContextFlow, compared with the aforementioned existing methods, where detailed discussions on related works are postponed to Appendix A.
In summary, our contributions are as follows:
  • We leverage the structure of local tissue organization and ligand–receptor communication to extract biologically meaningful features from spatial omics data and encode them into an informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
  • We design two novel integration schemes—cost-based and entropy-based—that incorporate prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
  • Comprehensive experiments on three regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline flow matching methods under both interpolation and extrapolation settings across multiple evaluation metrics that capture biological plausibility and statistical fidelity (Section 4).
2 Preliminaries
2.1 Flow Matching Basics
Flow matching (Lipman et al., 2023) is a simulation-free and sample-efficient generative framework for training continuous normalizing flows (Chen et al., 2018). Given a pair of source and target data distributions over $\mathbb{R}^{d}$ with densities $q_{0}=q(\bm{x}_{0})$ and $q_{1}=q(\bm{x}_{1})$, the problem task is to learn a time-varying velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, whose continuous evolution is captured by a function in the form of a neural-net-based model with weights $\theta$, that can transform $q_{0}$ to $q_{1}$ through integration via an ordinary differential equation (ODE). To be more specific, flow matching (FM) seeks to optimize $\theta$ by minimizing a simple regression loss between $u_{\theta}$ and a target time-varying velocity vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as follows:
Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani and others, 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator. From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$. The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$. While the flow matching objective in Equation 1 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
2.2 Conditional Flow Matching
The central idea of conditional flow matching is to express the target probability path via a mixture of more manageable conditional probability paths (Lipman et al., 2023). By marginalizing over a conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
$p_{t}(\bm{x})$ $=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},$ (2)
$u_{t}(\bm{x})$ $=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$
where $q(\bm{z})$ stands for the distribution of the conditioning variable $\bm{z}$, and the conditional probability path $p_{t}(\bm{x}|\bm{z})$ is constructed such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$. Theorem 1 of Lipman et al. (2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$.
To avoid the intractable integrals, (Lipman et al., 2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original FM objective in terms of gradient computation:
A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields, defined as follows:
$p_{t}(\bm{x}|\bm{z})$ $=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),$ (4)
$u_{t}(\bm{x}|\bm{z})$ $=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{x}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (5)
where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ stands for the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ denotes its derivative.
In particular, Lipman et al. (2023) selected $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$. Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{0},\mathbf{I})$ to a Gaussian distribution with mean $\bm{x}_{1}$ and covariance $\sigma^{2}\mathbf{I}$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2}\mathbf{I})$ for any target point $\bm{x}_{1}$. By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be validated. Tong et al. (2024) further generalized the application scope to arbitrary source distributions by setting:
$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\sigma_{t}(\bm{z})=\sigma.$ (6)
Note that the choice in Equation 6 satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$. According to Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
2.3 Flow Matching with Optimal Transport Couplings
The conditionals construction specified by Equation 6 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$. The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani and others, 2008). Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
Since the classical Kantorovich’s formulation has computational complexity that is cubic in terms of the sample size (see Appendix B.1 for its formal definition), a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
${\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}$ $\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})$ (7)
$\qquad\qquad+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$. For simplicity, we use $\pi^{*}_{\mathrm{eot}}(\epsilon)$ to denote the optimal solution to the minimization problem in Equation 7. Note that this optimization problem can also be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013). Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 14 in Appendix B.1 for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
3 Regularizing the Flow with Spatial Priors
3.1 Problem Formulation
We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data. Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points.For any $i\in\{1,2,\ldots,m+1\}$, let $q_{i}$ be the data distribution over $\mathbb{R}^{d}$ at time point $t_{i}$. Given $\{\mathbf{X}_{t_{i}}\}_{i\in\{1,2,\ldots,m+1\}}$, where $\mathbf{X}_{t_{i}}=\{\bm{x}_{i}(k)\}_{k=1}^{n_{i}}$ is the gene expressions at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$. This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points. Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, the conditional velocity field $u_{t}(\bm{x}|\bm{z})$ and its associated probability path $p(\bm{x}|\bm{z})$ are defined as:
$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),$
$u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (8)
where the conditioning variable $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 7. It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$. To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and to integrate biological prior knowledge inferred from other associated data modalities. Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension. However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 4(a) in Appendix G.1 for an illustration).
3.2 Spatial Priors and Transitional Plausibility
To faithfully model the context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the plausibility of transitions between locations and cell states across time points.
Spatial Smoothness. Tissues are well-organized systems. Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affect their states similarly and result in local smoothness in cell-type-specific expression. Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale. However, the same heterogeneity allows us to consider the spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2025) as a proxy for relative cell localization, which cannot change significantly across short time intervals. Therefore the aggregate expression within the microenvironment of each cell can be used to quantify the transitional plausibility in consecutive time points.
$\mathrm{SS}(c_{i},c_{j})$ $=\bigg\|\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)$ (9)
$\qquad\qquad-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\bigg\|_{2}^{2},$
where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ stands for the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile with respect to cell $c$.
Cell-Cell Communication Patterns. Cell–cell communication (CCC) has a critical role in the regulation of numerous biological processes, including development, apoptosis, and the maintenance of homeostasis in health and disease (Armingol et al., 2024). A major type of CCC is ligand–receptor (LR) signaling, in which ligands expressed by one cell bind to cognate receptors on another, initiating intracellular cascades that ultimately affect the state of the cell (i.e., its expression profile) (Armingol et al., 2021). There are numerous databases of prior knowledge of ligand-receptor binding and computational methods that use these databases to systematically link gene expression with the activity of ligand-receptor-mediated communication.
Specifically, we represent each cell $c_{i}$ by a feature vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair. The transitional plausibility $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 8 for an illustration). Formally, we define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
@@ 3.3 Flow Matching with Context-Aware OT Couplings
Transitional Plausibility Matrix. To begin with, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points. Specifically, denote by $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}$ $\left(c_{i}(k),c_{i+1}(l)\right)$ (11)
$+$ $(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right).$
@@
Prior-Regularized OT Couplings. The transitional plausibility matrices capture our spatially informed prior over cell-cell transitions between consecutive time points, which can naturally be incorporated into the EOT formulation (Equation 7) to promote couplings that maintain the structural and functional properties of tissue organization. We propose two techniques for prior integration:
Prior-Aware Cost Matrix (PACM). Consider the empirical counterpart of Equation 7 with respect to time $t_{i}$ and time $t_{i+1}$. Our first PACM approach incorporates the transitional plausibility matrix directly into the transport cost:
@@
where the $(k,l)$-th entry of the new prior-aware cost matrix is $\mathbf{C}_{i,i+1}]_{kl}=\alpha\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)[\mathbf{M}_{i,i+1}]_{kl}$, the transport plan $\Pi$ needs to satisfy the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility. If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 12 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$. This observation aligns with our assumption that such transitions should be more biologically implausible.
@@
Prior-Aware Entropy Regularization (PAER). While the previous PACM approach penalizes couplings according to our spatial priors, it defines a different OT problem with a modified cost function. Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds. Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison. This normalization, however, may result in couplings that deviate from their original counterparts (see Proposition 1 and Corollary 1 in Appendix D). Besides, selecting an appropriate $\alpha$ in Equation 12 introduces an additional layer of tuning, increasing computational overhead. Therefore, we propose the second PAER approach to integrate the biological priors without introducing additional hyperparameters:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ (13)
$\qquad\qquad-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\big(\log\big(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl}\big)-1\big)}_{\text{Prior-Aware Entropy Regularization}},$
@@
where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\dfrac{\exp(-[\mathbf{M}_{i,i+1}]_{kl})}{\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})}$ denotes the $(k,l)$-th entry of the prior joint probability matrix induced by the originally constructed $\mathbf{M}_{i,i+1}$. Note that the lower the cost $\mathbf{M}_{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}_{i,i+1}]_{kl}$, suggesting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$. The entropy regularization term in Equation 13 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
Table 2: Interpolation at the middle holdout time point for the Brain Regeneration dataset.
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
CTF-C 11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
CTF-C 11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
@@
ContextFlow. We apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problems defined in Equations 12 and 13 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 3.1.
The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings, is presented in Algorithm 1 in Appendix E.1. We also provide time complexity analysis in Appendix E.2. In particular, to apply the Sinkhorn algorithm to solve our PAER problem in Equation 13, we make use of the following theorem, a generalized result of Peyré et al. (2019).
@@
Theorem 3.1, proven in Appendix C, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices. When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 14. When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with EOT defined by Equation 7. This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow. By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.

### v3

@@ 4.2 Flow Matching with Context-Aware OT Couplings
Transitional Plausibility Matrix. To begin with, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points. Specifically, denote by $\mathbf{M}_{i}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
$\big[\mathbf{M}_{i}\big]_{kl}=\lambda\cdot\mathrm{SS}$ $\left(c_{i}(k),c_{i+1}(l)\right)+(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right).$ (6)
@@
Prior-Regularized OT Couplings. The transitional plausibility matrix encodes our spatially informed prior over cell-cell transitions between consecutive time points, which can naturally be incorporated into the EOT formulation (Equation 2) to promote couplings that maintain the structural and functional properties of tissue organization. In particular, we propose two techniques for prior integration:
Prior-Aware Cost Matrix (PACM). Consider $(t_{i},t_{i+1})$ and the empirical counterpart of Equation 2. Our first approach incorporates the transitional plausibility matrix directly into the transport cost:
@@
where the $(k,l)$-th entry $\mathbf{C}_{i,i+1}]_{kl}=\alpha\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)[\mathbf{M}_{i,i+1}]_{kl}$, the transport plan $\mathbf{\Pi}$ needs to satisfy the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility. If $\mathbf{M}_{i}]_{kl}$ is high, Equation 7 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$. This observation aligns with our assumption that such transitions are less biologically plausible.
@@
Prior-Aware Entropy Regularization (PAER). While PACM penalizes couplings according to our spatial priors, it defines a different OT problem with a modified cost function. Thus, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds. Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison. This normalization, however, may result in couplings that deviate from their original counterparts (see Theorem C.1 in Appendix C for a formal statement). Besides, selecting an appropriate $\alpha$ in Equation 7 introduces an additional layer of tuning, increasing computational overhead. To address the above issues, we propose the following PAER approach:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}-\epsilon\sum_{k,l}\Pi_{kl}\big(\log\big(\Pi_{kl}/[\widehat{\mathbf{M}}_{i}]_{kl}\big)-1\big),$ (8)
@@
where $\widehat{\mathbf{M}}_{i}]_{kl}=\exp(-[\mathbf{M}_{i}]_{kl})/{\sum_{l}\exp(-[\mathbf{M}_{i}]_{kl})}$ denotes the $(k,l)$-th entry of the prior joint probability matrix induced by $\mathbf{M}_{i}$. Note that the lower the cost $\mathbf{M}_{i}]_{kl}$, the higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$. Consequently, the proposed entropy regularization term in Equation 8 biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i}$ rather than a uniform baseline, providing a soft mechanism for incorporating the intended biological prior knowledge.
@@
ContextFlow. Finally, we apply the Sinkhorn algorithm (Cuturi, 2013) to obtain the spatial context-aware EOT couplings by solving optimization problems 7 and 8 for minibatch samples, and train the neural velocity vector field $u_{\theta}$ by minimizing the regression loss defined by Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z}_{i})$ and $u_{t}(\bm{x}|\bm{z}_{i})$ defined according to Equation 1. The pseudocode of ContextFlow is presented in Algorithm 1 in Appendix E.1, and we analyze its time complexity in Appendix E.3. In particular, to apply the Sinkhorn algorithm to solve our PAER problem, we make use of the following theorem, a generalization of Peyré et al. (2019).
@@
Theorem 4.1, proven in Appendix D.1, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost matrix $\mathbf{C}$ and the prior joint probability matrix $\mathbf{M}$. When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}\rightarrow\mathbf{\Pi}_{\mathrm{OT}}^{*}$, suggesting it recovers the standard OT couplings (Equation 15). When $\epsilon\rightarrow\infty$, $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, suggesting a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with EOT defined by Equation 2. This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow. By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.